Urban Drapers has a sister company named Super Carpeting Inc. (SCI). SCI just paid a dividend (Do) of $2.88 per share, and itplease answer last three questions:)
Answer #1

Current price = 2.88(1.06)/9% =$33.92

Dividend yield = expected dividend/current price

= 2.88(1.06)/33.92

= 9.00%

Stock price one year from today = 2.88(1.06)^2/(15%-6%)

=$35.96

Capital gains yield = (Price after one year- current price)/current price

=(35.96-33.92)/33.926

= 6.01%

I.e. 6%